OpenStudy (anonymous):

which fraction falls between 3/10 and 3/5 ?

OpenStudy (gorv):

3/20=0.15 3/5=0.6 so there arte lot of fraction betweem 0.15 and 0.60 as @Luis_Rivera said

hero (hero):

She hasn't listed the fractions yet

OpenStudy (anonymous):

Lol your right ! i didnt complete the question they give me the options of 5/10, 8/10 4/5 2/10

OpenStudy (anonymous):

you said between 0.15 and 0.60 I dont understand how any of the options fall in between ?

hero (hero):

Well, 5/10 = .50 8/10 = .80 4/5 = .80 2/10 = .20

OpenStudy (anonymous):

correction the question asked 3/10 between 3/5 lol ? sorry

OpenStudy (anonymous):

so 3/10 = 0.3 3/5 = 0.6 so the answer will be umm..2/10 ?

hero (hero):

No, it isn't. Sorry.

OpenStudy (anonymous):

grr um ok soo how do i get the answer lol

hero (hero):

2/10 = 0.2 right?

hero (hero):

If you created a number line: 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8... What falls between 0.3 and 0.6?

OpenStudy (anonymous):

0.4 and 0.5 so my answer will be 5/10 ?

hero (hero):

Yes..see how easy that was? You were making it more complicated

OpenStudy (anonymous):

haha yea i really was i really dont like fractions !! but yaay thank you so much :)

OpenStudy (gorv):

@Hero sorry i did not put zero before the point 3/20= 0.15 now it is correct thanks mate